OpenStudy (anonymous):

I've been told this problem is already in general form - If it is, what should I do with it? "Rewrite y = x2 + 14x + 29 in general form."

OpenStudy (anonymous):

leave it as it is

OpenStudy (anonymous):

Maybe they are telling you, that complete the square.

OpenStudy (anonymous):

yea its not in general form u have to move the y and complete the square twice

OpenStudy (anonymous):

If is that, y=x²+14x+29 To complete the square woul be sum both sides(14/2)² y+49=x²+14x+29+49 Factorizing y+49-29=(x+7)² y+20=(x+7)² Maybe is that.

OpenStudy (preetha):

I think the general form is the quadratic, ax2+bc+c=0